Gabrielle is 9 years older than Mikhail. The sum of their ages is 77. What is Mikhail's age?
Answer by TimothyLamb(4379)

x = Gabrielle age
y = Mikhail age
---
x + y = 77
x = y + 9
---
put the system of linear equations into standard form
---
x + y = 77
x - y = 9
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 43
y = 34
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---